# Wei Jiang    Jiayu Yang    Yongqi Zhai    Feng Gao    Ronggang Wang

### v9

@@ Abstract
Recently, learned image compression has achieved impressive performance. The entropy model, which estimates the distribution of the latent representation, plays a crucial role in enhancing rate-distortion performance. The latent representation contains channel-wise, local spatial, and global spatial correlations. However, existing global context modules rely on computationally intensive quadratic complexity computations to capture global correlations. The quadratic complexity imposes limitations on the potential of high-resolution image coding. Moreover, effectively capturing local, global, and channel-wise contexts with acceptable even linear complexity within a single entropy model remains a challenge. To address these limitations, we propose the Linear Complexity Multi-Reference Entropy Model (${}^{++}$). ${}^{++}$ effectively captures the diverse range of correlations inherent in the latent representation. Specifically, the latent representation is first divided into multiple slices. When compressing a particular slice, the previously compressed slices serve as its channel-wise contexts. To capture local contexts without sacrificing performance, we introduce a novel checkerboard attention module with linear complexity. Additionally, to capture global contexts, we propose the linear complexity attention-based global correlations capturing by leveraging the decomposition of the softmax operation. The attention map of the previously decoded slice is implicitly computed and employed to predict global correlations in the current slice. Based on ${}^{++}$, we propose image compression model ${}^{++}$. Extensive experimental evaluations demonstrate that our ${}^{++}$ achieves state-of-the-art performance, reducing BD-rate by $13.39\%$ on the Kodak dataset compared to VTM-17.0 in Peak Signal-to-Noise Ratio (PSNR). Furthermore, ${}^{++}$ exhibits linear GPU memory consumption with resolution, making it highly suitable for high-resolution image coding. Code and pre-trained models are available at https://github.com/JiangWeibeta/MLIC.
@@
wei.jiang1999@outlook.com, jiangwei@stu.pku.edu.cn
rgwang@pkusz.edu.cn
@@ 1 Introduction
Figure 1: BD-Rate-GPU Memory Consumption during inference on CLIC Professional Valid (Toderici et al., 2020) with 2K resolution. Our ${}^{++}$ achieves a better trade-off between performance and GPU memory consumption.
@@
Recently, various learned image compression models  (Wu et al., 2022; Guo et al., 2022; Ballé et al., 2017; Theis et al., 2017; Ballé et al., 2018; Minnen et al., 2018; Hu et al., 2020; Ma et al., 2020) have emerged, showcasing impressive performance gains. Notably, certain learned image compression models (Minnen & Singh, 2020; Cheng et al., 2020; Zou et al., 2022; Qian et al., 2020; Xie et al., 2021; He et al., 2022; Jiang et al., 2023b; Chen et al., 2021; Gao et al., 2021; Chen et al., 2022; Koyuncu et al., 2022; Duan et al., 2023; Liu et al., 2023; Jiang et al., 2023a; Fu et al., 2023a; Tang et al., 2023; Chen & Ma, 2023) are already comparable to the advanced traditional method VVC. These models predominantly rely on auto-encoders or variational auto-encoders (Kingma & Welling, 2014), and follow a process that involves transform, quantization, entropy coding, and inverse transform. Entropy coding plays an important role in boosting model performance. An entropy model is utilized to estimate the entropy of the latent representation. A powerful and accurate entropy model usually leads to fewer bits. Expanding contexts of entropy model in learned codecs plays the same role as expanding prediction modes in traditional codecs.
State-of-the-art learned image compression models (Cheng et al., 2020; Minnen & Singh, 2020; Xie et al., 2021; Zhu et al., 2022b; He et al., 2022; Jiang et al., 2023b, a) commonly enhance the entropy model by incorporating a hyper-prior module (Ballé et al., 2018) or a context module (Minnen et al., 2018). These additional modules enable the estimation of conditional entropy and the utilization of conditional probabilities for entropy coding. Context modules usually model probabilities and correlations in different dimensions, including local spatial context module, global spatial context module, and channel-wise context module. However, the current global context modules rely on computationally intensive quadratic complexity computations, which consume huge GPU memories and have slower encoding and decoding speed, imposing limitations on the potential for high-resolution image coding. Furthermore, effectively capturing local, global, and channel-wise contexts with acceptable even linear complexity within a single entropy model remains a challenge. To overcome the aforementioned limitations, we propose a novel linear complexity multi-reference entropy model. This entropy model effectively captures local spatial, global spatial, and channel-wise contexts with linear complexity and can be employed for efficient high-resolution image coding, which is denoted as ${}^{++}$ to differ from our prior work (Jiang et al., 2023b) presented at ACMMM 2023. Based on ${}^{++}$, we introduce ${}^{++}$, which achieve state-of-the-art performance.
@@
  • In order to exploit the global correlations within the latent representations, we divide them into slices and utilize the attention map of the previously decoded slice to predict the global correlations in the current slice. To overcome the quadratic complexity typically associated with the softmax operation in vanilla attention, we decompose it into two independent softmax operations, thereby achieving linear complexity without sacrificing performance. Additionally, we explore the global correlations from previous slices.
  • We design linear complexity multi-reference entropy model ${}^{++}$ which captures local spatial, global spatial and channel contexts, as well as hyper-prior side information. Based on ${}^{++}$, we propose ${}^{++}$, which achieves state-of-the-art performance. The complexity and GPU memory consumption of our model is linear with the resolution. Our proposed ${}^{++}$ achieved a better trade-off between complexity and performance as depicted in Fig. 1 .
In comparison to our previous work presented at ACMMM 2023 (Jiang et al., 2023b), ${}^{++}$ introduces several significant advancements. The primary distinction lies in the utilization of the proposed linear complexity global spatial context modules without sacrificing performance, as opposed to the quadratic complexity observed in our prior work. This achievement is primarily attributed to the division of the softmax operation, which eliminates the need for a specific order of tensor computation. Our proposed modules incorporate advanced techniques such as learnable position embedding and depth-wise residual bottlenecks (Jiang et al., 2023a). Furthermore, ${}^{++}$ captures inter-slice global correlations from all previous slices, in contrast to our previous work (Jiang et al., 2023b), which only considers correlations within the previous one slice. ${}^{++}$ exhibits several advantages, including reduced GPU memory consumption and faster encoding and decoding speed. The linear complexity makes our ${}^{++}$ highly suitable for high-resolution image coding. These advancements in ${}^{++}$ contribute to the field of image compression by offering improved efficiency and performance, while maintaining high-quality compression capabilities.
@@ 2.1 Learned Image Compression
Learned image compression (Ballé et al., 2017; Theis et al., 2017) aims to optimize the trade-off between distortion $\mathcal{D}$ and entropy, where entropy is typically measured in terms of bit-rate $\mathcal{R}$. Large bit-rate usually leads to lower distortion . Lagrange multiplier $\lambda$ is employed to adjust the weight of distortion to control the target bit-rate. The optimization target is
@@
where $\boldsymbol{x}$ represents the input image, $g_{a}$ transform the $\boldsymbol{x}$ to compact latent representation $\boldsymbol{y}$. $\boldsymbol{y}$ is quantized to $\hat{\boldsymbol{y}}$ for entropy coding. $\hat{\boldsymbol{x}}$ represents the decompressed image. $\theta$ and $\phi$ are parameters of $g_{a}$ and $g_{s}$. Since quantization is non-differentiable, it can be addressed during training by either adding uniform noise $\mathcal{U}(-0.5,0.5)$ (Ballé et al., 2017, 2018) or using the straight-through estimator (STE) (Theis et al., 2017). In particular, when uniform noise is added, the rate-distortion optimization target in Equation 1 is equivalent to evidence lower bound (ELBO) optimization in variational auto-encoders (Kingma & Welling, 2014). To enhance non-linearity, Generalized Divisive Normalization (GDN) (Ballé et al., 2015) layers or its variants (Qian et al., 2020) are employed. Additionally, self-attention (Vaswani et al., 2017; Liu et al., 2021; Lu et al., 2022; Zou et al., 2022; Liu et al., 2023; Guo et al., 2022), ensemble techniques (Wang et al., 2021), and block partition (Wu et al., 2022) are utilized in transform modules for more compact latent representations. In the basic model, a factorized or a non-adaptive density entropy model is adopted.
In subsequent works, a hyper-prior module (Ballé et al., 2018) is introduced to extract side information $\hat{\boldsymbol{z}}$ from $\boldsymbol{y}$. The hyper-prior model estimates the distribution of $\hat{\boldsymbol{y}}$ from $\hat{\boldsymbol{z}}$. A univariate Gaussian distribution is commonly employed for the hyper-prior. Some works extend it to a mean and scale Gaussian distribution (Minnen et al., 2018), asymmetric Gaussian distribution (Cui et al., 2021), Gaussian mixture model (Cheng et al., 2020; Liu et al., 2020b), and Gaussian-Laplacian-Logistic mixture model (Fu et al., 2023b) for more flexible distribution modeling.
@@ 2.2 Context-based Entropy Modeling
On the other hand, some approaches focus on modeling correlations between distant symbols. In (Qian et al., 2020), neighboring left and top symbols serve as bases for computing the similarity between the target symbol and its previous symbols. Guo et al. (Guo et al., 2022) employ the $L2$ distances of symbols to predict global casual dependencies among symbols. In (Kim et al., 2022), the side information is divided into global side information and local side information, introducing additional bits. However, these global context modules are typically combined with serial auto-regressive context modules, which further increase decoding latency. Moreover, existing global context modules (Qian et al., 2020; Guo et al., 2022; Jiang et al., 2023b) often exhibit quadratic complexity, making them challenging to apply in high-resolution image coding. Alternatively, they rely on extra side information (Kim et al., 2022), which increases the bit-rate.
@@
While some local and channel-wise context modules (Ma et al., 2021; He et al., 2022; Jiang et al., 2023b) have demonstrated impressive performance, effectively capturing local, global, and channel-wise contexts with acceptable even linear complexity within a single entropy model remains a challenge. Addressing these correlations has the potential to further enhance the performance of image compression models.
Figure 2: Visualization of channels of latent representation of Kodim19 extracted by Cheng’20 (Cheng et al., 2020) (optimized for MSE, $\lambda=0.0483$) to illustrate channel-wise redundancy.
Figure 3: Heatmap of spatial cosine similarity of latent representation of Kodim19 extracted by Cheng’20 (Cheng et al., 2020) (optimized for MSE, $\lambda=0.0483$) to visualize global spatial and local spatial redundancy.
Notations Explanation
$\boldsymbol{x},\hat{\boldsymbol{x}}$ Input and decoded image
$\boldsymbol{y},\hat{\boldsymbol{y}}$ Non-quantized and quantized latent representation
$\hat{\boldsymbol{y}}^{i}$ The $i$-th slice of $\hat{\boldsymbol{y}}$
$\hat{\boldsymbol{y}}_{ac},\hat{\boldsymbol{y}}_{na}$ Anchor and non-anchor part of $\hat{\boldsymbol{y}}$
$\boldsymbol{z},\hat{\boldsymbol{z}}$ Non-quantized and quantized side information
$g_{ep}$ Entropy parameter module
$\mu,\sigma$ Mean and scale of $\hat{\boldsymbol{y}}$
$g_{a},g_{s}$ Analysis and synthesis transform
$h_{a},h_{s}$ Hyper analysis and synthesis
$g_{ch}$ Channel-wise context module
$g_{lc,ckbd}$ Vanilla checkerboard context module
$g_{lc,attn}$ Shifted Window-based Checkerboard Attention
$g_{gc,intra}$ Intra-slice global spatial context module
$g_{gc,inter}$ Inter-slice global spatial context module
${\boldsymbol{\Phi}}_{h},\boldsymbol{\Phi}_{ch},{\boldsymbol{\Phi}}_{lc}$ Hyper-prior, channel-wise and local spatial context
${\boldsymbol{\Phi}}_{gc,intra}$ Intra-slice global spatial context
${\boldsymbol{\Phi}}_{gc,inter}$ Inter-slice global spatial context
MEM ${}^{++}$ Linear complexity
multi-reference entropy model
$M,N,S$ Channel number of $\boldsymbol{y}$, $\boldsymbol{z}$, and $\hat{\boldsymbol{y}}^{i}$
$K$ Kernel size of local spatial context module
Q, AE, AD Quantization, arithmetic encoding and decoding
@@
Figure 4: The overall architecture of ${}^{++}$. $\downarrow$ means down-sampling. $\uparrow$ means up-sampling. / means stride equals $1$. Red line is the dataflow during decoding. ${\boldsymbol{x}}$ is the input image and $\hat{\boldsymbol{x}}$ is the reconstructed image. $\boldsymbol{y}$ is the latent representation and $\hat{\boldsymbol{y}}$ is the quantized latent representation. $\hat{\boldsymbol{y}}^{i}$ is the $i$-th slice of $\hat{\boldsymbol{y}}$. .

### v10

# MLIC++: Linear Complexity Multi-Reference Entropy Modeling for Learned Image Compression

@@ Abstract
The latent representation in learned image compression encompasses channel-wise, local spatial, and global spatial correlations, which are essential for the entropy model to capture for conditional entropy minimization. Efficiently capturing these contexts within a single entropy model, especially in high-resolution image coding, presents a challenge due to the computational complexity of existing global context modules. To address this challenge, we propose the Linear Complexity Multi-Reference Entropy Model (MEM++). Specifically, the latent representation is partitioned into multiple slices. For channel-wise contexts, previously compressed slices serve as the context for compressing a particular slice. For local contexts, we introduce a shifted-window-based checkerboard attention module. This module ensures linear complexity without sacrificing performance. For global contexts, we propose a linear complexity attention mechanism. It captures global correlations by decomposing the softmax operation, enabling the implicit computation of attention maps from previously decoded slices. Using MEM++ as the entropy model, we develop the image compression method MLIC++. Extensive experimental results demonstrate that MLIC++ achieves state-of-the-art performance, reducing BD-rate by $13.39\%$ on the Kodak dataset compared to VTM-17.0 in Peak Signal-to-Noise Ratio (PSNR). Furthermore, MLIC++ exhibits linear computational complexity and memory consumption with resolution, making it highly suitable for high-resolution image coding. Code and pre-trained models are available at https://github.com/JiangWeibeta/MLIC.
@@
wei.jiang1999@outlook.com, rgwang@pkusz.edu.cn
@@ 1 Introduction
Figure 1: Left: BD-Rate-GPU Memory Consumption during inference on CLIC Professional Valid (Toderici et al., 2020) with 2K resolution. Our MLIC++ achieves a better trade-off between performance and GPU memory consumption. Right: Reconstruction comparison on “vita-vilcina-3055” from CLIC Professional Valid (Toderici et al., 2020) dataset. The reconstruction of MLIC++ has the best visual quality.
@@
Recently, various learned image compression models  (Wu et al., 2022; Guo et al., 2022; Ballé et al., 2017; Theis et al., 2017; Ballé et al., 2018; Minnen et al., 2018; Hu et al., 2020; Ma et al., 2020) have emerged, showcasing impressive performance gains. Notably, certain learned image compression models (Minnen & Singh, 2020; Cheng et al., 2020; Zou et al., 2022; Qian et al., 2020; Xie et al., 2021; He et al., 2022; Jiang et al., 2023; Chen et al., 2021; Gao et al., 2021; Chen et al., 2022; Koyuncu et al., 2022; Duan et al., 2023b; Liu et al., 2023; Jiang et al., 2024; Duan et al., 2023a) are already comparable to the advanced traditional method VVC. These models predominantly rely on auto-encoders or variational auto-encoders (Kingma & Welling, 2013), and follow a process that involves transform, quantization, entropy coding, and inverse transform. Entropy coding plays an important role in boosting model performance. An entropy model is utilized to estimate the entropy of the latent representation. A powerful and accurate entropy model usually leads to fewer bits. Expanding contexts of entropy model in learned codecs plays the same role as expanding prediction modes in traditional codecs.
State-of-the-art learned image compression models (Cheng et al., 2020; Minnen & Singh, 2020; Xie et al., 2021; Zhu et al., 2022b; He et al., 2022; Jiang et al., 2023, 2024; Lin et al., 2023) commonly enhance the entropy model by incorporating a hyper-prior module (Ballé et al., 2018) or a context module (Minnen et al., 2018). These additional modules enable the estimation of conditional entropy and the utilization of conditional probabilities for entropy coding. Context modules usually model probabilities and correlations in different dimensions, including local spatial context module, global spatial context module, and channel-wise context module. However, the current global context modules rely on computationally intensive quadratic complexity computations, which consume huge GPU memories and have slower encoding and decoding speed, imposing limitations on the potential for high-resolution image coding. Furthermore, effectively capturing local, global, and channel-wise contexts with acceptable even linear complexity within a single entropy model remains a challenge. To overcome the aforementioned limitations, we propose a novel linear complexity multi-reference entropy model. This entropy model effectively captures local spatial, global spatial, and channel-wise contexts with linear complexity and can be employed for efficient high-resolution image coding, which is denoted as MEM++ to differ from our prior work (Jiang et al., 2023) presented at ACMMM 2023. Based on MEM++, we introduce MLIC++, which achieve state-of-the-art performance.
@@
  • For enhanced context modeling efficiency, we decompose the softmax operations in vanilla attention into two independent softmax operations. This reduces computational complexity to linear without compromising performance. Additionally, we explore the global correlations
  • We design linear complexity We design linear complexity multi-reference entropy model MEM++ which captures local spatial, global spatial and channel contexts, as well as hyper-prior side information. Using MEM++ as the entropy model, we develop MLIC++, which achieves state-of-the-art performance with linear complexity. Our proposed MLIC++ achieved a better trade-off between complexity and performance as depicted in Fig. 1.
In comparison to our previous work presented at ACMMM 2023 (Jiang et al., 2023), MLIC++ introduces several significant advancements. The primary distinction lies in the utilization of the proposed linear complexity global spatial context modules without sacrificing performance, as opposed to the quadratic complexity observed in our prior work. This achievement is primarily attributed to the division of the softmax operation, which eliminates the need for a specific order of tensor computation. Our proposed modules incorporate advanced techniques such as learnable position embedding and depth-wise residual bottlenecks (Jiang et al., 2024). Furthermore, MLIC++ captures inter-slice global correlations from all previous slices, in contrast to our previous work (Jiang et al., 2023), which only considers correlations within the previous one slice. MLIC++ exhibits several advantages, including reduced GPU memory consumption and faster encoding and decoding speed. The linear complexity makes our MLIC++ highly suitable for high-resolution image coding. To demonstrate the superiority of MLIC++, we conduct comprehensive performance and complexity comparisons against existing methods (Cheng et al., 2020; Minnen & Singh, 2020; Qian et al., 2020; Xie et al., 2021; Qian et al., 2022; Zhu et al., 2022b; Wang et al., 2022; Zhu et al., 2022a; Zou et al., 2022; He et al., 2022; Liu et al., 2023) across multiple datasets (Toderici et al., 2020; Kodak, 1993; Asuni & Giachetti, 2014) and resolutions (Liu et al., 2020a), extending beyond the preliminary experiments presented in the workshop version (Jiang & Wang, 2023). These advancements in MLIC++ contribute to the field of image compression by offering improved efficiency and performance, while maintaining high-quality compression capabilities.
@@ 2.1 Learned Image Compression
Learned image compression (Ballé et al., 2017; Theis et al., 2017) aims to optimize the trade-off between distortion $\mathcal{D}$ and entropy, where entropy is typically measured in terms of bit-rate (Theis et al., 2017; Yang et al., 2023; Ballé et al., 2017, 2018, 2020) $\mathcal{R}$. Large bit-rate usually leads to lower distortion . Lagrange multiplier $\lambda$ is employed to adjust the weight of distortion to control the target bit-rate. The optimization target is
@@
where $\boldsymbol{x}$ represents the input image, $g_{a}$ transform the $\boldsymbol{x}$ to compact latent representation $\boldsymbol{y}$. $\boldsymbol{y}$ is quantized to $\hat{\boldsymbol{y}}$ for entropy coding. $\hat{\boldsymbol{x}}$ represents the decompressed image. $\theta$ and $\phi$ are parameters of $g_{a}$ and $g_{s}$. Since quantization is non-differentiable, it can be addressed during training by either adding uniform noise $\mathcal{U}(-0.5,0.5)$ (Ballé et al., 2017, 2018) or using the straight-through estimator (STE) (Theis et al., 2017). In particular, when uniform noise is added, the rate-distortion optimization target in Equation 1 is equivalent to evidence lower bound (ELBO) optimization in variational auto-encoders (Kingma & Welling, 2013). To enhance non-linearity, Generalized Divisive Normalization (GDN) (Ballé et al., 2015) layers or its variants (Qian et al., 2020) are employed. Additionally, self-attention (Vaswani et al., 2017; Liu et al., 2021; Lu et al., 2022; Zou et al., 2022; Liu et al., 2023; Guo et al., 2022), ensemble techniques (Wang et al., 2021), and block partition (Wu et al., 2022) are utilized in transform modules for more compact latent representations. In the basic model, a factorized or a non-adaptive density entropy model is adopted.
In subsequent works, a hyper-prior module (Ballé et al., 2018) is introduced to extract side information $\hat{\boldsymbol{z}}$ from $\boldsymbol{y}$. The hyper-prior model estimates the distribution of $\hat{\boldsymbol{y}}$ from $\hat{\boldsymbol{z}}$. A univariate Gaussian distribution is commonly employed for the hyper-prior. Some works extend it to a mean-scale Gaussian distribution (Minnen et al., 2018), asymmetric Gaussian distribution (Cui et al., 2021), Gaussian mixture model (Cheng et al., 2020; Liu et al., 2020b), and Gaussian-Laplacian-Logistic mixture model (Fu et al., 2023) for more flexible distribution modeling.
@@ 2.2 Context-based Entropy Modeling
On the other hand, some approaches focus on modeling correlations between distant symbols. In (Qian et al., 2020), neighboring left and top symbols serve as bases for computing the similarity between the target symbol and its previous symbols. Guo et al. (Guo et al., 2022) employ the $L2$ distances of symbols to predict global casual dependencies among symbols. In (Kim et al., 2022), the side information is divided into global side information and local side information, introducing additional bits. However, these global context modules are typically combined with serial auto-regressive context modules, which further increase decoding latency. Moreover, existing global context modules (Qian et al., 2020; Guo et al., 2022; Jiang et al., 2023) often exhibit quadratic complexity, making them challenging to apply in high-resolution image coding. Alternatively, they rely on extra side information (Kim et al., 2022), which increases the bit-rate.
@@
While some local and channel-wise context modules (Ma et al., 2021; He et al., 2022; Jiang et al., 2023) have demonstrated impressive performance, effectively capturing local, global, and channel-wise contexts with acceptable even linear complexity within a single entropy model remains a challenge. Addressing these correlations has the potential to further enhance the performance of image compression models.
Figure 2: Visualization of channels of latent representation of Kodim19 extracted by Cheng’20 (Cheng et al., 2020) (optimized for MSE, $\lambda=0.0483$) to illustrate channel-wise redundancy. These channels are nearest-neighbor upsampled for visualization.
Figure 3: Heatmap of spatial cosine similarity of latent representation of Kodim19 extracted by Cheng’20 (Cheng et al., 2020) (optimized for MSE, $\lambda=0.0483$) to visualize global spatial and local spatial redundancy. The heatmap is nearest-neighbor upsampled for visualization.
Notations Explanation Notations Explanation
$\boldsymbol{x},\hat{\boldsymbol{x}}$ Input and decoded image $\boldsymbol{y},\hat{\boldsymbol{y}}$ Non-quantized and quantized latent representation
$\hat{\boldsymbol{y}}^{i}$ The $i$-th slice of $\hat{\boldsymbol{y}}$ $\boldsymbol{z},\hat{\boldsymbol{z}}$ Non-quantized and quantized side information
$\hat{\boldsymbol{y}}_{ac},\hat{\boldsymbol{y}}_{na}$ Anchor and non-anchor part of $\hat{\boldsymbol{y}}$ $g_{ep}$ Entropy parameter module
$\mu,\sigma$ Mean and scale of $\hat{\boldsymbol{y}}$ $g_{a},g_{s}$ Analysis and synthesis transform
$h_{a},h_{s}$ Hyper analysis and synthesis $g_{ch}$ Channel-wise context module
$g_{lc,ckbd}$ Vanilla checkerboard context module $g_{lc,attn}$ Shifted Window-based Checkerboard Attention
$g_{gc,intra}$ Intra-slice global spatial context module $g_{gc,inter}$ Inter-slice global spatial context module
${\boldsymbol{\Phi}}_{h},\boldsymbol{\Phi}_{ch},{\boldsymbol{\Phi}}_{lc}$ Hyper-prior, channel-wise and local spatial context ${\boldsymbol{\Phi}}_{gc,intra}$ Intra-slice global spatial context
${\boldsymbol{\Phi}}_{gc,inter}$ Inter-slice global spatial context MEM ++ Linear complexity multi-reference entropy model
$M,N,S$ Channel number of $\boldsymbol{y}$, $\boldsymbol{z}$, and $\hat{\boldsymbol{y}}^{i}$ $K$ Kernel size of local spatial context module
@@
$N$ $M$ $S$ $K$ Entropy Model
$192$ $320$ $32$ $5$ MEM++($g_{lc,attn},g_{ch},g_{gc,intra},g_{gc,inter}$)
Table 2: Settings of MLIC++ and MEM++.